A hand driven tire pump has a piston with a 2.1cm diameter and a maximum stroke of 32 cm.

a) How much work do you do in one stroke if the average gauge pressure is 2.9 x 10^5 N/m^2?

b) What average force do you exert on the piston, neglecting friction and gravitational force?
Answer #1

a)

W = PΔV

ΔV = πd2/4 /l =   π * 0.021^2 /4 *0.32= 0.000110835389 m3

W = 2.9*10^5*0.000110835389 =32.1422628 J

so work done =32.1422628 J

B)

x = displacement in one stroke =0.32 m

W =F*x

F =W/x = 32.1422628/0.32 =100.444571 N answer

****************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..